# Entire Large Solutions for Competitive Semilinear Elliptic Systems with General Nonlinearities Satisfying Keller–Osserman Conditions

### v1

@@ 3 Proof of the Theorems
which has a (maximal) $C^{2}$ solution $y$ on $\left[R_{0},R_{a}\right)$ by standard ODE theory.
Define the Keller–Osserman primitive and we denote
$\mathcal{C}_{f}(s):=\int_{s_{\ast}}^{s}\frac{d\tau}{\ell_{f}(\tau)},\qquad s>s_{\ast},$
where $s_{\ast}>0$ is chosen so that $\ell_{f}(\tau)>0$ on $s_{\ast},\infty)$. Then $\mathcal{C}_{f}$ is strictly increasing on $(s_{\ast},\infty)$, with
$\mathcal{C}_{f}^{\prime}(s)=\frac{1}{\ell_{f}(s)},\qquad\mathcal{C}_{f}^{\prime\prime}(s)=-\frac{\ell_{f}^{\prime}(s)}{\ell_{f}(s)^{2}}\leq 0,$
and, by the Keller–Osserman condition, $\mathcal{C}_{f}(\infty):=\lim_{s\rightarrow\infty}\mathcal{C}_{f}(s)<\infty$.
Set $W(r):=\mathcal{C}_{f}\bigl{(}y(r)\bigr{)}$. Differentiating and using (3.6) yields
$W^{\prime}(r)=\frac{y^{\prime}(r)}{\ell_{f}(y(r))},\qquad\bigl{(}r^{n-1}W^{\prime}(r)\bigr{)}^{\prime}=\frac{r^{n-1}}{\ell_{f}(y)}\Bigl{(}y^{\prime\prime}+\frac{n-1}{r}y^{\prime}\Bigr{)}-\frac{r^{n-1}\ell_{f}^{\prime}(y)}{\ell_{f}(y)^{2}}\,y^{\prime 2}\;\leq\;r^{n-1}p(r).$
Integrating from $R_{0}$ to $r$ and using $W^{\prime}(R_{0})=y^{\prime}(R_{0})/\ell_{f}(a)=0$ gives
@@
Dividing by $r^{n-1}$ and integrating once more from $R_{0}$ to $r$ leads to the Keller–Osserman estimate
$\mathcal{C}_{f}\bigl{(}y(r)\bigr{)}\;\leq\;\mathcal{C}_{f}(a)\;+\;Q_{R_{0}}(r),\qquad Q_{R_{0}}(r):=\int_{R_{0}}^{r}t^{1-n}\!\left(\int_{R_{0}}^{t}s^{n-1}p(s)\,ds\right)\,dt.$ (3.7)
@@
so $Q_{R_{0}}(\infty)<\infty$, and moreover $Q_{R_{0}}(\infty)\downarrow 0$ as $R_{0}\uparrow\infty$.
Choosing $R_{0}$ large so that $Q_{R_{0}}(\infty)$ is small, and $a$ close enough to $u(R_{0})$ but still smaller than $u(R_{0})$, we can ensure
$\mathcal{C}_{f}(a)+Q_{R_{0}}(\infty)>\mathcal{C}_{f}(\infty),$
which forces blow-up of $y$ at some finite $R_{a}>R_{0}$.
Now compare $u$ and $y$ on $A_{R_{0},R}$ for $R<R_{a}$. On $\partial A_{R_{0},R}$ we have $u(R_{0})>a=y(R_{0})$ and, for $R$ close to $R_{a}$, $y(R)$ arbitrarily large while $u$ is bounded. By the weak maximum principle applied to uyu-y in $\{u<y\}$ (note the reversed inequality at $R_{0}$), we reach a contradiction with the boundedness of $u$ (as in Step 1 of the proof of Theorem 2.2). Therefore $u$ must be unbounded, hence $u(r)\rightarrow+\infty$ as $r\rightarrow\infty$. ∎
@@
Proof. The argument is symmetric, using $\ell_{g}$, $q$, and a barrier $w$ with $w(R_{1})<v(R_{1})$ to obtain the same type of contradiction. ∎
Conclusion. Combining the two claims yields, under (2.3),
@@
Step 7: Conclusion. We have constructed a positive radial entire solution $(u,v)$ of (1.1), established
$W\left(r\right)=u\left(r\right)+v\left(r\right)\rightarrow\infty\text{ as }r\rightarrow\infty,$
proved the quantitative lower bound (2.2), and shown that at least one component or both diverges under (1.2)-(2.3). This completes the proof of Theorem 2.3. ∎

### v2

@@ 3 Proof of the Theorems
which has a maximal $C^{2}$ solution $y$ on $R_{0},R_{a})$ by standard ODE theory.
Define the Keller–Osserman primitive
$\mathcal{C}_{f}(s):=\int_{s_{\ast}}^{s}\frac{d\tau}{\ell_{f}(\tau)},\quad s>s_{\ast},$
where $s_{\ast}>0$ is chosen so that
$\ell_{f}(\tau)>0\text{ on }[s_{\ast},\infty).$
Then $\mathcal{C}_{f}$ is strictly increasing on $(s_{\ast},\infty)$, with
$\mathcal{C}_{f}^{\prime}(s)=\frac{1}{\ell_{f}(s)},\quad\mathcal{C}_{f}^{\prime\prime}(s)=-\frac{\ell_{f}^{\prime}(s)}{\ell_{f}(s)^{2}}\leq 0,$
and, by the Keller–Osserman condition, $\mathcal{C}_{f}(\infty)<\infty$.
Set $W(r):=\mathcal{C}_{f}\big(y(r)\big)$. Differentiating and using (3.7) yields
$W^{\prime}(r)=\frac{y^{\prime}(r)}{\ell_{f}(y(r))},\quad\big(r^{n-1}W^{\prime}(r)\big)^{\prime}=\frac{r^{n-1}}{\ell_{f}(y)}\left(y^{\prime\prime}+\frac{n-1}{r}y^{\prime}\right)-\frac{r^{n-1}\ell_{f}^{\prime}(y)}{\ell_{f}(y)^{2}}\,(y^{\prime})^{2}\leq r^{n-1}p(r).$
Integrating from $R_{0}$ to $r$ and using $W^{\prime}(R_{0})=y^{\prime}\left(R_{0}\right)/\mathcal{C}_{f}(a)=0$ gives
@@
Dividing by $r^{n-1}$ and integrating again from $R_{0}$ to $r$ yields the Keller–Osserman estimate
$\mathcal{C}_{f}\big(y(r)\big)\leq\mathcal{C}_{f}(a)+Q_{R_{0}}(r),\quad Q_{R_{0}}(r):=\int_{R_{0}}^{r}t^{1-n}\left(\int_{R_{0}}^{t}s^{n-1}p(s)\,ds\right)dt.$ (3.8)
@@
so $Q_{R_{0}}(\infty)<\infty$ and $Q_{R_{0}}(\infty)\downarrow 0$ as $R_{0}\uparrow\infty$.
Blow-up of the barrier. Since $\mathcal{C}_{f}(\infty)<\infty$, we can first choose $R_{0}$ large so that $Q_{R_{0}}(\infty)$ is as small as desired, then choose $a$ close enough to $u(R_{0})$ but still smaller than $u(R_{0})$, so that $\mathcal{C}_{f}(a)\;+\;Q_{R_{0}}(\infty)\;>\;\mathcal{C}_{f}(\infty)$. Because $Q_{R_{0}}$ is continuous and strictly increasing, there exists a finite $R_{a}\in(R_{0},\infty)$ with
$\mathcal{C}_{f}(a)\;+\;Q_{R_{0}}(R_{a})\;=\;\mathcal{C}_{f}(\infty).$ (3.9)
By (3.8), this forces $\mathcal{C}_{f}\bigl(y(r)\bigr)\nearrow\mathcal{C}_{f}(\infty)$ as $r\uparrow R_{a}$, hence $\lim_{r\uparrow R_{a}}y(r)\;=\;+\infty$, i.e., $y$ is a (finite-radius) blow-up supersolution for (3.6) on $R_{0},R_{a})$.
Maximum principle comparison and contradiction. Fix any $R\in(R_{0},R_{a})$. Since $u$ is assumed bounded on $R_{0},\infty)$ while $y(r)\rightarrow\infty$ as $r\uparrow R_{a}$, we may choose $R$ sufficiently close to $R_{a}$ so that
$y(R)>\sup_{r\geq R_{0}}u(r).$
In particular,
$u(R_{0})-y(R_{0})=u(R_{0})-a>0,\qquad u(R)-y(R)<0.$
. enDefine w:=uyw:=u-y on the annulus
$A_{R_{0},R}:=\bigl\{x\in\mathbb{R}^{n}:\;R_{0}<r=|x|<R\bigr\}\,\}.$
Using (3.6) and (3.7), together with the monotonicity of $l_{f}$, we obtain on the set where $u\geq y$:
$\Delta w=\Delta u-\Delta y>p(r)\,l_{f}(u)-p(r)\,l_{f}(y)\geq 0.$
Let
$r_{\ast}:=\inf\{\,r\in(R_{0},R]:w(r)\leq 0\,\}.$
By the boundary values above, $r_{\ast}\in(R_{0},R)$ and
$w(r)>0\quad\text{for }r\in[R_{0},r_{\ast}),\qquad w(r_{\ast})=0.$
Thus $w$ is nonnegative on $R_{0},r_{\ast}]$ and strictly positive inside, with $w$ attaining its minimum 0 at the boundary point $r_{\ast}$ of the domain $(R_{0},r_{\ast})$. Moreover, on $(R_{0},r_{\ast})$ we have $u\geq y$ and hence $\Delta w>0$.
The situation is illustrated in Figure, where the blue curve denotes $u$, the red curve the barrier $y$.
[…]
Figure 3: Schematic of the first contact $r_{\ast}$ where $u$ (blue) and $y$ (red) intersect: solid for $r<r_{\ast}$, dashed for $r>r_{\ast}$. On $(R_{0},r_{\ast})$ we have w=uy>0w=u-y>0 and $\Delta w>0$, so the Hopf boundary lemma yields $w^{\prime}(r_{\ast})>0$, contradicting the first-contact requirement $w^{\prime}(r_{\ast})\leq 0$.
By the (radial) weak maximum principle applied to $w$ on $(R_{0},r_{\ast})$, $w$ cannot achieve a strict interior minimum; by the Hopf boundary point lemma at $r_{\ast}$, the outward normal derivative must satisfy
$w^{\prime}(r_{\ast})>0\text{ (outward normal points to ward increasing }r\text{)}.$
However, $r_{\ast}$ is the first contact point where $w$ crosses from positive to nonpositive values, so necessarily $w^{\prime}(r_{\ast})\leq 0$. This contradiction shows that such $r_{\ast}$ cannot exist. The next Figure shows $u$ vs. $y$ and why the “forbidden crossing” at $r_{\ast}$ can’t happen.
[…]
Figure 4: $u=y$ separates the regions $u>y$ and $u<y$. The trajectory starts in $u>y$ at $R_{0}$ and touches the diagonal for the first time at $r_{\ast}$. The maximum principle forbids crossing into $u<y$ here.
Hence w(r)=u(r)y(r)>0w(r)=u(r)-y(r)>0 on $R_{0},R)$. Letting $R\uparrow R_{a}$ yields $u(r)\geq y(r)$ for all $r\in[R_{0},R_{a})$. Since $y(r)\rightarrow\infty$ as $r\uparrow R_{a}$, we must have $u(r)\rightarrow\infty$ as $r\uparrow R_{a}$, which contradicts the assumed boundedness of $u$ (and contradicts the prescribed initial value comparison $u(R_{0})>a$ together with the smoothness of $u$ on $R_{0},\infty)$). Therefore, the assumption that $u$ is bounded while $v\rightarrow\infty$ is impossible.
@@
Proof of Claim 2. The argument is symmetric, using $\ell_{g}$, $q$, and a barrier $w$ with $w(R_{1})<v(R_{1})$ to obtain the same type of contradiction. ∎∎
Step 7: Conclusion.
Proof.
We have now completed the construction of a positive, radial, entire solution
$(u,v)\in C^{2}([0,\infty))\times C^{2}([0,\infty))$
to the system (1.1) under the standing assumptions (H1)–(H3). From the previous steps, we have established that the sum
$W(r):=u(r)+v(r)$
satisfies
$W(r)\longrightarrow+\infty\quad asr\rightarrow\infty,$
and moreover obeys the quantitative lower bound (2.2) derived via the Keller–Osserman transform.
We have also shown that:
  • Without any additional conditions added to (H1)-(H3), at least one of the two components $u$ or $v$ must diverge to $\infty$ as $r\rightarrow\infty$.
  • Under the supplementary Keller–Osserman condition at infinity (2.3), both components necessarily diverge:
@@
Thus, the constructed pair $(u,v)$ is a positive, radial, entire solution of (1.1) with the stated asymptotic properties, completing the proof of Theorem 2.3. ∎